# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

### v2

@@ 4.3. Results
Dependence on $K$. In distributed computing, scaling up typically results in proportional speed improvements. In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup. This unpredictability is starkly illustrated with LeNet-5 and VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies. Figures 7, and 8 (left) demonstrate that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication. These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources. However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (left part of Figures 910), the expected behavior starts to emerge. Especially for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies. Communication cost, however, increases with $K$ for all methods except Synchronous, which maintains constant communication irrespective of worker count, but at the expense of orders of magnitude higher communication overhead. Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers. Our FDA variants consistently outperform FedAdam, FedAvgM, and Synchronous in communication efficiency, as demonstrated across all experiments in Figures 5-7. Specifically, they require up to 30 times less communication than FedAdam, 4 times less than FedAvgM, and up to 2.5 orders of magnitude less than Synchronous.
@@
FDA: Dependence on $\Theta$. The variance threshold $\Theta$ can be seen as a lever in balancing communication and computation; essentially, it calibrates the trade-off between these two costs. A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence. This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 7-10). Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 9 and 10. It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
@@
FDA: Choice of $\Theta$. The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 7,  89, and 10. Therefore, identifying this range becomes crucial. To this end, we conducted extensive exploratory testing to estimate the $\Theta$ ranges for each learning task which are predominantly influenced by the number of parameters $d$ of the DNN. Within this context, $\Theta$ values outside the desirable range exhibit notable effects: below this range, the training process mimics Synchronous or Local-SGD approaches with small $\tau$, while exceeding it leads to non-convergence. Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 2), thereby investigating different computation and communication trade-offs. For instance, in the ARIS-HPC environment (https://www.hpc.grnet.gr/en/hardware-2/) with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by the computation cost, rendering communication concerns negligible. In such contexts, lower $\Theta$ values are favored due to their computational efficiency. On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
To assist researchers in selecting the variance threshold, Figure 11 presents empirical estimations for $\Theta$ across three distinct learning settings: (1) FL (assuming a common channel of $0.5$Gbps), (2) Balanced communication-computation equilibrium setting, and (3) our HPC environment at the ARIS supercomputer.
$\Theta_{FL}$ $=4.91\cdot 10^{-5}\cdot d$
$\Theta_{B}$ $=3.89\cdot 10^{-5}\cdot d$
$\Theta_{HPC}$ $=2.74\cdot 10^{-5}\cdot d$
FDA: Linear vs. Sketch. In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 234, and 5. This suggests that the precision of the variance approximation is not critical. However, in all experiments within the more intricate transfer learning scenario, LinearFDA requires approximately $1.5$ times more communication than SketchFDA to fine-tune the deep ConvNeXtLarge model to equivalent performance levels (Figure 12). In light of these findings, we conclude the following: for straightforward and less demanding tasks, LinearFDA is the recommended option due to its simplicity and lower complexity per local state computation. On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, if communication-efficiency is paramount.
5. Conclusions
In this paper, we introduced Federated Dynamic Averaging (FDA), an innovative, adaptive and communication-efficient algorithm for distributed deep learning. Essentially, FDA makes informed, dynamic decisions on when to terminate training rounds based on approximations of the model variance. Through extensive experiments across diverse datasets and learning tasks, we demonstrated that FDA significantly reduces communication overhead (often by orders of magnitude) without a corresponding increase in computation or compromise in model performance — contrary to the typical trade-offs encountered in the literature. Furthermore, we showed that FDA is robust to data heterogeneity and inherently mitigates overfitting. Our results push the limits of modern communication-efficient distributed deep learning, paving the way for scalable, dynamic, and broadly applicable strategies.
@@ Acknowledgements.
This work was supported by the EU project CREXDATA under Horizon Europe agreement No. 101092749. Moreover, this work was supported by computational time granted from the National Infrastructures for Research and Technology S.A. (GRNET S.A.) in the National HPC facility - ARIS - under project ID pa230902-fda1.

### v4

@@ 4.3. Results
Dependence on $K$. In distributed computing, scaling up typically results in proportional speed improvements. In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup (Yu et al., 2019). This unpredictability is starkly illustrated with LeNet-5 and VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies. Figures 8, and 9 (top) demonstrate that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication. These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources. However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (top part of Figures 1011), the expected behavior starts to emerge. Especially for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies. Communication cost, however, increases with $K$ for all methods except Synchronous, which maintains constant communication irrespective of worker count, but at the expense of orders of magnitude higher communication overhead. Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers. Our FDA variants consistently outperform FedAdam, FedAvgM, and Synchronous in communication efficiency, as demonstrated across all experiments in Figures 8-11. Specifically, they require up to 30 times less communication than FedAdam, 4 times less than FedAvgM, and up to 2.5 orders of magnitude less than Synchronous.
@@
FDA: Dependence on $\Theta$. The variance threshold $\Theta$ can be seen as a lever in balancing communication and computation; essentially, it calibrates the trade-off between these two costs. A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence. This impact of $\Theta$ is consistently observed across both FDA strategies, and all learning tasks, and data heterogeneity settings (Figures 8-11). Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 10 and 11. It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
FDA: Choice of $\Theta$. The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 8,  910, and 11. Therefore, identifying this range becomes crucial. To this end, we conducted extensive exploratory testing to estimate the $\Theta$ ranges for each learning task which are predominantly influenced by the number of parameters $d$ of the DNN. Within this context, $\Theta$ values outside the desirable range exhibit notable effects: below this range, the training process mimics Synchronous or Local-SGD approaches with small $\tau$, while exceeding it leads to non-convergence. Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 2), thereby investigating different computation and communication trade-offs. For instance, in the ARIS-HPC environment with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by the computation cost, rendering communication concerns negligible. In such contexts, lower $\Theta$ values are favored due to their computational efficiency. On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
@@
To assist researchers in selecting the variance threshold, Figure 12 presents empirical estimations for $\Theta$ across three distinct learning settings:
  1. (1)
    FL, assuming a common channel of $0.5$Gbps, where
    $\Theta_{FL}$ $=4.91\cdot 10^{-5}\cdot d$
  2. (2)
    Balanced communication-computation equilibrium, where
    $\Theta_{B}$ $=3.89\cdot 10^{-5}\cdot d$
  3. (3)
    Our HPC environment at the ARIS supercomputer, where
    $\Theta_{HPC}$ $=2.74\cdot 10^{-5}\cdot d$
FDA: Linear vs. Sketch. In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 346, and 6. This suggests that the precision of the variance approximation is not critical. However, in all experiments within the more intricate transfer learning scenario, LinearFDA requires approximately $1.5$ times more communication than SketchFDA to fine-tune the deep ConvNeXtLarge model to equivalent performance levels (Figure 13). In light of these findings, we conclude the following: for straightforward and less demanding tasks, LinearFDA is the recommended option due to its simplicity and lower complexity per local state computation. On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, if communication-efficiency is paramount.
@@
5. Conclusions and Future Work
In this paper, we introduced Federated Dynamic Averaging (FDA), an innovative, adaptive and communication-efficient algorithm for distributed deep learning. Essentially, FDA makes informed, dynamic decisions on when to synchronize the local models based on approximations of the model variance. Through extensive experiments across diverse datasets and learning tasks, we demonstrated that FDA significantly reduces communication overhead (often by orders of magnitude) without a corresponding increase in computation or compromise in model performance—contrary to the typical trade-offs encountered in the literature. Furthermore, we showed that FDA is robust to data heterogeneity and inherently mitigates over-fitting. Our results push the limits of modern communication-efficient distributed deep learning, paving the way for more scalable, dynamic, and broadly applicable strategies.
An interesting direction for future work is whether the value of $\Theta$ can be dynamically adjusted in order to achieve (or not to exceed) a target average bandwidth consumption. Since the expected behavior is that the communication cost decreases when $\Theta$ increases, such an approach seems feasible (i.e., increasing $\Theta$ when the bandwidth consumption is higher than what is desired), especially by using statistics. We plan to look into this extension in the future.
@@ Acknowledgements.
We wish to thank the anonymous reviewers and the meta-reviewer for their insightful comments and suggestions. This work was supported by the EU project CREXDATA under Horizon Europe agreement No. 101092749. Moreover, this work was supported by computational time granted from the National Infrastructures for Research and Technology S.A. (GRNET S.A.) in the National HPC facility - ARIS - under project ID pa230902-fda1.